A 15-kg rock is dropped from rest on the earth and reaches the ground in 1.75 s. When it is


dropped from the same height on Saturn’s satellite Enceladus, it reaches the ground in 18.6 s.


(a) What is the acceleration due to gravity on Enceladus? (b) How long will it take for the


rock to reach the ground on Earth if it was released 100 m above the Earth’s surface? (c) in


Enceladus?

1
Expert's answer
2022-04-21T13:02:23-0400

a)

"h=\frac{g_0 t_0^2}2=15~m,"

"g=\frac{2h}{t^2}=0.087~\frac m{s^2},"

b)

"t_1=\sqrt{\frac{2h'}{g_0}}=4.5~s,"

c)

"t_2=\sqrt{\frac{2h'}{g}}=48~s."